OpenStudy (anonymous):

Solve, using the substitution method. y – 2x = 8 16 + 4x = 2y A. The only solution is (24, 0). B. There are an infinite number of solutions. C. The only solution is (1, 10). D. There is no solution.

jimthompson5910 (jim_thompson5910):

Start with y - 2x = 8

jimthompson5910 (jim_thompson5910):

do you see how you would isolate y in y - 2x = 8 ?

OpenStudy (anonymous):

yah, so it would be: y=2x+8? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

once you know that y = 2x+8, you can replace every copy of y in the second equation with 2x+8

jimthompson5910 (jim_thompson5910):

16 + 4x = 2y 16 + 4x = 2( y ) 16 + 4x = 2( 2x+8 ) ... replace every copy of y with 2x+8 now solve for x

OpenStudy (anonymous):

k so now you need to distribute it, so 2*2x=4x and 2*8= 16 right?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

so 16 + 4x = 4x + 8 now we need to isolate the x

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

wait so now it would be 16 + 4x - 8/4 = x i think i did it wrong..

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

not sure how you're getting that

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

oh btw, I just noticed this, but you distributed incorrectly

jimthompson5910 (jim_thompson5910):

2( 2x+8 ) should be 4x + 16

OpenStudy (anonymous):

oh yah, i guess i forgot to multiply the 2*8

jimthompson5910 (jim_thompson5910):

so you really have 16 + 4x = 4x + 16

jimthompson5910 (jim_thompson5910):

If you rearrange terms, you go from 16 + 4x = 4x + 16 to 4x + 16 = 4x + 16

jimthompson5910 (jim_thompson5910):

and we have something in the form x = x, which is ALWAYS true for any real number x

OpenStudy (anonymous):

so it would be B. ?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

k thnx!!!!